9:30a – 12:20a
- Restart analysis for project 2
- Probe making: check PCRs, continue if good, restart if not.
- launch STORM analysis for last few days of imaging
- see notes.
- Test PCRs
- Clean, reorganize, recover gel bench. Lots of empty tip boxes, empty buffers,
- set up T7 reactions. Run overnight
- PCR gel: (D12, F11, F12, G1, G2, G3, G5, G6)
- running D10 and D08 data on Monet.
- New storm-analysis repository version not working on Tuck / Cajal.
- wrote to Hazen for suggestions in getting storm-analysis to work correctly.
- Need to modify .pth file inside Python27 folder. Except we don’t want everyone modifying this file to be different.
- Fixed startup to add python paths for windows.
SetPythonPaths =['set PYTHONPATH=%PYTHONPATH%;',stormAnalysisPath,'; && '];
STORM imaging plans
- Bogdan will image tonight and Saturday night.
- I will image Friday and Sunday night.
- Change of plans, I will image on both STORM2 and STORM4 tonight.
- confirm E05 fails. F12b also works quite well (have F12 already from last night).
- E03 small probe set surprisingly seems to work. Setting up to image E03 on STORM2 overnight
- image beads on STORM4, setting up to test D12rna on STORM4
- running E11 analysis on Tuck and Cajal
- Running F08 analysis on Cajal
- D12rna, spots highly variable. Not totally convinced this worked, I thought I had a much higher frequency of spots per cell last time I did D12. Actually, I don’t appear to have recorded a D12 rna dataset yet. I have multiple E10s. And a D12rna negative control. Should do the E10 rna neg control and the D10 rna soon.
- Background very high on D12 rna. 47C 1hr washout not sufficient. (Need the 60C oven). Try 60C washout tomorrow. Current background is borderline for imaging.
- very provactive: http://www.plosbiology.org/article/info%3Adoi%2F10.1371%2Fjournal.pbio.1000419
- seems to me active regions
- I don’t know about 9 nm separation distances of non-point source objects.